# On the Mechanisms of Collaborative Learning in VAE Recommenders

### v1

@@ 4 Experiments
  • First, we assess benefits of the proposed personalized item alignment approach compared to vanilla VAE-based CF (Multi-VAE (liang2018variational)) on these benchmark datasets.
@@ 4.1 Effectiveness of Personalized Item Alignment
Table 1 presents the performance of our framework, which adds personalized item alignment to Multi-VAE, and RecVAE (shenbin2020recvae) on the MovieLens-20M, Netflix and Million Song datasets respectively. We follow the preprocessing procedure from (liang2018variational). The detailed data preprocessing steps and train/validation/test split methodology are presented in Section D.1. Our code for reproducibility is available at https://github.com/amazon-science/PIAVAE.
@@
On the basis of offline results, we run one week of A/B testing in September 2025 for the Multi-VAE + PIA algorithm on one streaming platform of Amazon. The approach was implemented as an offline system, with weekly training considering a 3-month window for collecting streaming behavior, and daily inference for active customers. The personalized scores computed daily include about $25$ millions of users and $4000$ movies. The performance of our system was compared with a statistical baseline (control group) on 2 movie cards present on the Homepage and on the Movie page. As shown in Table 3, our approach outperformed the control group significantly, with improved performance on the card click rate by $117\%-267\%$ (per daily view) and $123\%-283\%$ (per daily user view). Since its launch, we have observed the performance of the ML system to be remarkably stable in playtime and click rate metrics.
@@ 4.2 Latent Space Visualization
We present t‑SNE visualizations (maaten2008visualizing) of the latent spaces learned under three settings: ① clean input, ② input masking, and ③ input masking with personalized item alignment, to examine the correspondence between the geometry of the input space and the latent space.
@@ 4.3 Performance of user groups with different number of interactions
Our proposed framework provides both a well-structured latent space and the capacity to capture global collaborative signals. As a result, we expect it to benefit users across groups, including cold-start, neutral, and warm-start users. In particular, cold-start performance 10.1145/3485447.3512048; 10.1145/3640457.3688125; 10.1145/3705328.3748109 is of particular importance in industry settings. To assess this, we partition the test-set users based on their number of interactions and evaluate the performance of our method within each group.
@@ 5 Conclusion
In this work, we analyzed how collaboration emerges in VAE–CF and showed that it is fundamentally governed by latent proximity: SGD updates are shared within a data‑dependent sharing radius, clean inputs bias the model toward local collaboration, and global signals can be induced by either the $\beta$–KL/prior pathway (near‑uniform contraction of latent distances, with collapse risk if over‑used) or by input masking (stochastic neighborhood mixing with potential drift). Guided by these insights, we introduced Personalized Item Alignment (PIA), a training‑only regularizer that attaches learnable item anchors and softly pulls masked encodings toward each user’s anchor centroid. PIA preserves instance information, stabilizes the geometry under masking, and promotes semantically grounded global mixing without adding test‑time overhead. Empirically, PIA improves over vanilla VAE–CF on standard benchmarks and in an A/B test on a large‑scale streaming platform, with ablations across user‑activity strata and latent‑space visualizations corroborating the theory.
THE USE OF LARGE LANGUAGE MODELS
We used a large language model (ChatGPT) to help with editing this paper. It was only used for simple tasks such as fixing typos, rephrasing sentences for clarity, and improving word choice. All ideas, experiments, and analyses were done by the authors, and the use of LLMs does not affect the reproducibility of our work.
We also used ChatGPT to assist with proof verification and theorem refinement. Our workflow involved providing initial drafts to ChatGPT, which would then suggest improvements to the mathematical presentation and formatting. We subsequently edited and refined these suggestions.
@@
  • The pseudo-code of the algorithm is provided in Section C.
  • Detailed experimental settings and implementation details are described in Section D.
@@ Appendix B Related Works
$\beta$-scheduling. Liang et al. (liang2018variational) introduce a $\beta$-scaling factor to modulate the strength of the regularization $\mathrm{KL}\!\left(q_{\phi}(\mathbf{z}\mid\mathbf{x})\,\|\,p(\mathbf{z})\right)$, while Long et al. (Long2019PreventingPC) propose gradually increasing this weight over training to mitigate posterior collapse.
Flexible priors and architectures. Several works replace the standard normal prior with richer alternatives to better match the data. Examples include VampPrior and its hierarchical variants (e.g., HVamp) (tomczak2018vae; kim2019enhancing), as well as implicit or learned priors (walker2022implicit). RecVAE (shenbin2020recvae) combines a redesigned encoder–decoder, a composite prior, input-dependent $\beta(\mathbf{x})$ rescaling, alternating training, and a non-denoising decoder. Other lines incorporate user-dependent priors (karamanolakis2018item) or impose an arbitrary target prior via adversarial training (zhang2018adversarial).
@@
Appendix C Algorithm
The model is trained by optimizing the objective defined in Eq. (8). However, for the hyperparameter $\lambda_{\text{alignment}}$, which controls the strength of the personalized item alignment regularization, instead of consider it as fixed hyper-parameter, we gradually increase its value during training .Specifically, we use the validation set to monitor whether the latent space is getting trapped in a local optimum i.e., when the validation performance does not improve after $\rho$ consecutive epochs. If such a case is detected, we increase $\lambda_{\text{alignment}}$ by a scaling factor $\lambda_{\text{scale}}$, as detailed in Algorithm 1.
Algorithm 1 Personalized Item Alignment VAE
1:Initialize: Models: encoder $q_{\phi}$, sequence of invertible transformations $f_{1},...,f_{T}$, decoder $p_{\theta}$ and item-embeddings $E$.Hyper-parameters: $\lambda_{\text{A}}$, $\lambda_{\text{scale}}$ and $\rho$.Variables: $\textit{best\_val\_epoch}:=0$, $\textit{best\_val\_ndcg}:=0$
2:for epoch in n_epochs do
3:  // Training
4:  for iter in iterations do
5:   Sample a mini-batch $\mathbf{x}$
6:   $\mathbf{z}_{0}\sim q_{\phi}(\cdot\mid\mathbf{x})$ // using reparametrization trick
7:   $\mathbf{z}_{T}=\left(f_{T}\circ...\circ f_{1}\right)\left(\mathbf{z}_{0}\right)$
8:   Update $q_{\phi},f_{1},...f_{T},p_{\theta}$ and $E$ based on $\mathcal{L}_{\text{PIA-VAE}}(\mathbf{x};\theta,\phi,f_{1},...,f_{T})$ in Equation (8)
9:  end for
10:  //Validation
11:  Compute nDCG@K on validation set: epoch_ndcg
12:  if $\textit{epoch\_ndcg}>\textit{best\_val\_ndcg}$ then
13:   $\textit{best\_val\_ndcg}:=\textit{epoch\_ndcg}$
14:   $\textit{best\_val\_epoch}:=\textit{epoch}$
15:  end if
16:  // Increase $\lambda_{\text{A}}$ if training is stuck in a local optimum, i.e., when the validation performance does not improve after $\rho$ consecutive epochs.
17:  if $\textit{best\_val\_epoch}<\textit{epoch}+\rho$ then
18:   $\lambda_{\text{A}}:=\lambda_{\text{scale}}\times\lambda_{\text{A}}$
19:  end if
20:end for
21:Return: the optimal encoder $q_{\phi}$ and decoder $p_{\theta}$ at best_val_epoch.
Appendix D Experimental Settings
D.1 Dataset
We validate our analysis using three real-world recommendation datasets: MovieLens-20M [^2] https://grouplens.org/datasets/movielens/20m/, Netflix [^3] https://www.kaggle.com/netflix-inc/netflix-prize-data and Million Song (MSD)  [^4] http://millionsongdataset.com, where each record consists of a user-item pair along with a rating that the user has given to the item. We follow the preprocessing procedure from MultVAE (liang2018variational). For MovieLens-20M and Netflix, we retain users who have rated at least five movies and treat ratings of four or higher as positive interactions. For MSD, we keep only users with at least 20 songs in their listening history and songs that have been listened to by at least 200 users.
The user data is split into training, validation, and test sets as presented in Table 5.
Table 5: Dataset statistics
@@ D.1 Dataset
For every user in the training set, we utilize all interaction history, whereas for users in the validation or test set, a fraction of the history (80%) is used to predict the remaining interaction.
D.2 Implementation Details
@@
We use a batch size of 500 and train the model for 200 epochs using the Adam optimizer with a learning rate of $1\times 10^{-3}$ across all experiments. The specific hyperparameters $\lambda_{\text{A}}$, $\rho$, and $\lambda_{\text{scale}}$ are selected based on validation performance. We set $\rho=5$, $\lambda_{\text{scale}}=2$ and $\lambda_{\text{A}}=8$ consistently across all experiments, as the model is not particularly sensitive to these values.
Implementation details and instructions for reproducing all experimental results and visualizations are provided in the accompanying source code.
Model Architecture
We adopt the same architecture as used in MultVAE liang2018variational for the base encoder, with additional layers for the deeper encoder and the encoder with normalizing flows. All encoders are followed by a shared decoder. The specific structures are as follows:
  • Base Encoder: A simple feedforward encoder with two linear layers:
    $\mathbb{R}^{|I|}\rightarrow\mathbb{R}^{600}\rightarrow\mathbb{R}^{200}$
  • Deeper Encoder: An extended version of the base encoder with two additional linear layers:
    $\mathbb{R}^{|I|}\rightarrow\mathbb{R}^{600}\rightarrow\mathbb{R}^{200}\underset{\text{Two additional layers}}{\underbrace{\rightarrow\mathbb{R}^{200}\rightarrow\mathbb{R}^{200}}}$
  • Encoder with Normalizing Flows: Enhances expressiveness by applying two Inverse Autoregressive Flow (IAF) layers kingma2016improved after the base encoder:
    $\mathbb{R}^{|I|}\rightarrow\mathbb{R}^{600}\rightarrow\mathbb{R}^{200}\underset{\text{Two additional layers}}{\underbrace{\rightarrow\text{IAF}_{1}\rightarrow\text{IAF}_{2}}}$
  • Decoder: Shared across all models, mapping the 200-dimensional latent representation back to the input space:
    $\mathbb{R}^{200}\rightarrow\mathbb{R}^{|I|}$
h! Model Recall@20 Recall@50 NDCG@100 Shallow Encoder 0.400 0.533 0.429 Shallow Encoder + Alignment 0.401 0.537 0.431 Uplift (%) 0.25 0.75 0.47 Deeper Encoder 0.400 0.538 0.427 Deeper Encoder + Alignment 0.403 0.540 0.430 Uplift (%) 0.75 0.37 0.70 Shallow (IAF) Encoder 0.401 0.540 0.431 Shallow (IAF) Encoder + Alignment 0.406 0.545 0.437 Uplift (%) 1.25 0.93 1.39
Table 6: Impact of Posterior Flexibility on Alignment Performance (MovieLens-20M). Shallow Encoder refers to a model with 2-layer encoders, Deeper-Encoder uses 4-layer encoders, and Shallow (IAF) Encoder combines a 2-layer encoder with 2 invertible transformations.
VAE models for collaborative filtering (CF) applications typically employ shallow network architectures as a means of alleviating the issue of posterior collapse. However, this limitation reduces the model’s representational capacity and restricts posterior flexibility. In this constrained setting, we find that PIA becomes less effective i.e., its regularization offers little performance gain. This is reflected by the nDCG results in Table 6.
When increasing the depth of the encoder, the benefits of PIA become more pronounced, as shown in Table 6 by comparing Deeper-Encoder with and without alignment. However, Deeper-Encoder models also exhibit a greater tendency toward posterior collapse Dai2019TheUS, which ultimately undermines overall recommendation quality (see Table 6, comparing Shallow-Encoder and Deeper-Encoder). These findings underscore a critical trade-off in VAE-based CF models: while increased posterior flexibility—achieved through deeper architectures—enhances the effectiveness of personalized item alignment, it also intensifies the risk of posterior collapse, which in turn can degrade recommendation performance. Conversely, shallow architectures mitigate collapse but yield inflexible posteriors, limiting the benefits of alignment. Thus, the utility of alignment is closely tied to the model’s capacity to learn expressive yet stable latent representations.
To address this issue, we propose to enhance flexible posteriors while avoiding posterior collapse by using normalizing flows (NF) tabak2013family; kingma2016improved; rezende2015variational with deeper architectures. Normalizing flows builds flexible posterior distributions utilizing a sequence of $T$ invertible transformations to transform the original posterior $q_{\phi}(\mathbf{z}|\mathbf{x})$ into a flexible posterior $q_{\phi}(\mathbf{z}_{T}|\mathbf{x})$ as follows,
$\mathbf{z}_{0}\sim q_{\phi}(\mathbf{z}_{0}|\mathbf{x}),\quad\mathbf{z}_{t}=\mathbf{f}_{t}(\mathbf{z}_{t-1})\quad\forall\>t=1,\ldots T,$ (11)
where $\mathbf{f}_{t}$’s are trainable, invertible-transformations, such as invertible linear-time transformation rezende2015variational or inverse autoregressive network kingma2016improved.
We then apply the KL regularization in the transformed latent space $\mathcal{Z}_{T}$, i.e., $\text{KL}\bigl(q_{\phi}(\mathbf{z}_{T}|\mathbf{x})\|p(\mathbf{z})\bigr)$, rather than in the original latent space $\mathcal{Z}_{0}$. This design enables the posterior in the original latent space $\mathcal{Z}_{0}$ to remain flexible and better capture variations across different inputs. Owing to the invertibility of the transformations, the latent variables in the transformed space $\mathcal{Z}_{T}$ retain identifiability, even under strong regularization. This leads to a more robust and structured latent representation in VAE-based CF when using normalizing flows as follows:
$\mathcal{L}_{\text{VAE-NF}}(\mathbf{x};\theta,\phi,f_{1},...,f_{T})=-\mathbb{E}_{\mathbf{z}_{T}\sim q_{\phi}(\mathbf{z}_{T}|\mathbf{x}_{u})}\bigl[\log p_{\theta}(\mathbf{x}|\mathbf{z}_{T})\bigr]+\beta D_{\text{KL}}\bigl(q_{\phi}(_{T}|\mathbf{x})\parallel p(\mathbf{z})\bigr)$ (12)
where the sampling $\mathbf{z}\sim q_{\phi}(\mathbf{z}_{T}|\mathbf{x}_{u})$ is performed as follows: $\mathbf{b}\sim\operatorname{Ber}(\bm{\alpha})\rightarrow\mathbf{x}_{h}=\mathbf{x}\odot\mathbf{b}\rightarrow\text{with }\bm{\mu},\bm{\sigma}^{2}=q_{\phi}(\mathbf{x}_{h}),\bm{\epsilon}\sim\mathcal{N}(\bm{0},(\bm{1})\rightarrow\mathbf{z}_{0}=\bm{\mu}+\bm{\epsilon}\bm{\sigma}\rightarrow\mathbf{z}_{T}=f_{T}\circ\circ\circ f_{1}(\mathbf{z}_{0})$ and the probability density function of the last transformation can be computed as:
$\log q_{\phi}(\mathbf{z}_{T}|\mathbf{x})=\log q_{\phi}(\mathbf{z}_{0}|\mathbf{x})-\sum_{t=1}^{T}\log\det\left|\frac{d\mathbf{z}_{t}}{d\mathbf{z}_{t-1}}\right|.$ (13)
The effectiveness of the Shallow (IAF) Encoder posterior is shown from Table 6, where Shallow (IAF) Encoder not only achieves stronger base performance but also demonstrates clear improvements when alignment is applied.
D.3 Baselines
@@
  • Matrix factorization (MF); we consider MF trained with ALS with uniform weights (hu2008collaborative), which is a simple and computationally efficient baseline, and also weighted matrix factorization (wMF) (hu2008collaborative);
  • Regularization based on item-item interactions; here we selected GRALS (rao2015collaborative) that employs graph regularization;
  • Linear models; we have chosen full-rank models EASE steck2019embarrassingly and a low-rank model PLRec (sedhain2016practical);
  • Nonlinear autoencoders; here we consider the shallow autoencoder CDAE (wu2016collaborative), variational autoencoder MultVAE (liang2018variational), and its successors: RaCT (lobel2019towards) and RecVAE (shenbin2020recvae).

### v2

@@ 4 Experiments
  • First, we assess benefits of the proposed personalized item alignment approach compared to vanilla VAE-based CF (Multi-VAE (Liang et al., 2018)) on these benchmark datasets.
@@ 4.1 Effectiveness of Personalized Item Alignment
Table 1 presents the performance of our framework, which adds personalized item alignment to Multi-VAE, and RecVAE (Shenbin et al., 2020) on the MovieLens-20M, Netflix and Million Song datasets respectively. We follow the preprocessing procedure from (Liang et al., 2018). The detailed data preprocessing steps and train/validation/test split methodology are presented in Section C.1. Our code for reproducibility is publicly available at https://github.com/amazon-science/PIAVAE.
@@
On the basis of offline results, we run one week of A/B testing in September 2025 for the Multi-VAE + PIA algorithm on one streaming platform of Amazon. The approach was implemented as an offline system, with weekly training considering a 3-month window for collecting streaming behavior, and daily inference for active customers. The personalized scores computed daily include about $25$ millions of users and $4000$ movies. The performance of our system was evaluated on 2 movie cards present on the Homepage and on the Movie page. 50% of the customers were exposed to the algorithm (treatment group) for the duration of the experiment, while the rest was exposed to the baseline algorithm (control group). As shown in Table 3, our approach outperformed the control group significantly, with improved performance on the card click rate by $117\%-267\%$ (per daily view) and $123\%-283\%$ (per daily user view). For statistical validation, we report the p-value which evaluates the mean difference between control and treatment groups, as well as the Bayesian probability that the mean difference is positive, considering a Normal-Normal conjugate historic prior which allows for closed form solutions to the posterior distribution. For click rate metrics we observed $\text{p-value}=0.000$ and $(\text{prob}>0)=1.000$; and for playtime, we observed $\text{p-value}=0.000$ and $(\text{prob}>0)=0.997$. Since its launch, we have observed the performance of the ML system to be remarkably stable in playtime and click rate metrics.
@@ 4.2 Latent Space Visualization
We present t‑SNE visualizations (Maaten and Hinton, 2008) of the latent spaces learned under three settings: ① clean input, ② input masking, and ③ input masking with personalized item alignment, to examine the correspondence between the geometry of the input space and the latent space.
@@ 4.3 Performance of user groups with different number of interactions
Our proposed framework provides both a well-structured latent space and the capacity to capture global collaborative signals. As a result, we expect it to benefit users across groups, including cold-start, neutral, and warm-start users. In particular, cold-start performance Xu et al. (2022); Monteil et al. (2024); Liang et al. (2025) is of particular importance in industry settings. To assess this, we partition the test-set users based on their number of interactions and evaluate the performance of our method within each group.
@@ 5 Conclusion
In this work, we analyzed how collaboration emerges in VAE–CF and showed that it is fundamentally governed by latent proximity: SGD updates are shared within a data‑dependent sharing radius, clean inputs bias the model toward local collaboration, and global signals can be induced by either the $\beta$-KL/prior pathway (near‑uniform contraction of latent distances, with collapse risk if over‑used) or by input masking (stochastic neighborhood mixing with potential drift). Guided by these insights, we introduced Personalized Item Alignment (PIA), a training‑only regularizer that attaches learnable item anchors and softly pulls masked encodings toward each user’s anchor centroid. PIA preserves instance information, stabilizes the geometry under masking, and promotes semantically grounded global mixing without adding test‑time overhead. Empirically, PIA improves over vanilla VAE–CF on standard benchmarks and in an A/B test on a large‑scale streaming platform, with ablations across user‑activity strata and latent‑space visualizations corroborating the theory.
Limitations.
The benefits of capturing global collaborative signals still depend heavily on how well the input masking is designed (most current research relies on Bernoulli masking). If the masking is too noisy, even with alignment mechanisms, the model may struggle to learn meaningful representations. Therefore, a promising direction for future research is to explore more effective masking strategies that better support global collaboration.
Acknowledgments
We would like to thank Vy Vo for the valuable discussions during the preparation of this paper.
@@ Acknowledgments
  • Detailed experimental settings and implementation details are described in Section C.
  • The pseudo-code of the algorithm is provided in Section C.2.
  • Additional experiments on parameter sensitivity analysis are presented in Section D.
@@ Appendix B Related Works
$\beta$-scheduling. Liang et al. (Liang et al., 2018) introduce a $\beta$-scaling factor to modulate the strength of the regularization $\mathrm{KL}\!\left(q_{\phi}(\mathbf{z}\mid\mathbf{x})\,\|\,p(\mathbf{z})\right)$, while Long et al. (Long et al., 2019) propose gradually increasing this weight over training to mitigate posterior collapse.
Flexible priors and architectures. Several works replace the standard normal prior with richer alternatives to better match the data. Examples include VampPrior and its hierarchical variants (e.g., HVamp) (Tomczak and Welling, 2018; Kim and Suh, 2019), as well as implicit or learned priors (Walker et al., 2022). RecVAE (Shenbin et al., 2020) combines a redesigned encoder–decoder, a composite prior, input-dependent $\beta(\mathbf{x})$ rescaling, alternating training, and a non-denoising decoder. Other lines incorporate user-dependent priors (Karamanolakis et al., 2018) or impose an arbitrary target prior via adversarial training (Zhang et al., 2018).
@@
Appendix C Experimental Settings
C.1 Dataset
We validate our analysis using three real-world recommendation datasets: MovieLens-20M [^2] https://grouplens.org/datasets/movielens/20m/, Netflix [^3] https://www.kaggle.com/netflix-inc/netflix-prize-data and Million Song (MSD)  [^4] http://millionsongdataset.com, where each record consists of a user-item pair along with a rating that the user has given to the item. We follow the preprocessing procedure from MultVAE (Liang et al., 2018). For MovieLens-20M and Netflix, we retain users who have rated at least five movies and treat ratings of four or higher as positive interactions. For MSD, we keep only users with at least 20 songs in their listening history and songs that have been listened to by at least 200 users.
Table 5: Dataset statistics.
@@ C.1 Dataset
The user data is split into training, validation, and test sets as presented in Table 5. For every user in the training set, we utilize all interaction history, whereas for users in the validation or test set, a fraction of the history (80%) is used to predict the remaining interaction.
C.2 Implementation Details
@@
We use a batch size of 500 and train the model for 200 epochs using the Adam optimizer with a learning rate of $1\times 10^{-3}$ across all experiments. The specific hyperparameters $\lambda_{\text{A}}$, $\rho$, and $\lambda_{\text{scale}}$ are selected based on validation performance. An ablation study on the sensitivity of $\rho$, $\lambda_{\text{scale}}$ and $\lambda_{\text{A}}$ with respect to model performance is provided in Section D.1.
Algorithm.
The model is trained by optimizing the objective defined in Eq. (8). However, for the hyperparameter $\lambda_{\text{A}}$, which controls the strength of the personalized item alignment regularization, instead of consider it as fixed hyper-parameter, we gradually increase its value during training .Specifically, we use the validation set to monitor whether the latent space is getting trapped in a local optimum i.e., when the validation performance does not improve after $\rho$ consecutive epochs. If such a case is detected, we increase $\lambda_{\text{A}}$ by a scaling factor $\lambda_{\text{scale}}$, as detailed in Algorithm 1.
Algorithm 1 Personalized Item Alignment VAE
1:Initialize: Models: encoder $q_{\phi}$, decoder $p_{\theta}$ and item-embeddings $E$.Hyper-parameters: $\lambda_{\text{A}}$, $\lambda_{\text{scale}}$ and $\rho$.Variables: $\textit{best\_val\_epoch}:=0$, $\textit{best\_val\_ndcg}:=0$
2:for epoch in n_epochs do
3:  // Training
4:  for iter in iterations do
5:   Sample a mini-batch $\mathbf{x}$
6:   $\mathbf{z}\sim q_{\phi}(\cdot\mid\mathbf{x})$ // using reparametrization trick
7:   Update $q_{\phi},p_{\theta}$ and $E$ based on $\mathcal{L}_{\text{PIA-VAE}}(\mathbf{x};\theta,\phi)$ in Equation (8)
8:  end for
9:  //Validation
10:  Compute nDCG@K on validation set: epoch_ndcg
11:  if $\textit{epoch\_ndcg}>\textit{best\_val\_ndcg}$ then
12:   $\textit{best\_val\_ndcg}:=\textit{epoch\_ndcg}$
13:   $\textit{best\_val\_epoch}:=\textit{epoch}$
14:  end if
15:  // Increase $\lambda_{\text{A}}$ if training is stuck in a local optimum, i.e., when the validation performance does not improve after $\rho$ consecutive epochs.
16:  if $\textit{best\_val\_epoch}<\textit{epoch}+\rho$ then
17:   $\lambda_{\text{A}}:=\lambda_{\text{scale}}\times\lambda_{\text{A}}$
18:  end if
19:end for
20:Return: the optimal encoder $q_{\phi}$ and decoder $p_{\theta}$ at best_val_epoch.
C.3 Baselines
@@
  • Matrix factorization (MF); we consider MF trained with ALS with uniform weights (Hu et al., 2008), which is a simple and computationally efficient baseline, and also weighted matrix factorization (wMF) (Hu et al., 2008);
  • Regularization based on item-item interactions; here we selected GRALS (Rao et al., 2015) that employs graph regularization;
  • Linear models; we have chosen full-rank models EASE Steck (2019) and a low-rank model PLRec (Sedhain et al., 2016);
  • Nonlinear autoencoders; here we consider the shallow autoencoder CDAE (Wu et al., 2016), variational autoencoder MultVAE (Liang et al., 2018), and its successors: RaCT (Lobel et al., 2019) and RecVAE (Shenbin et al., 2020).
Appendix D Additional Experiments
D.1 Parameter sensitivity analysis on $\lambda_{\text{A}}$ $\rho$ and $\lambda_{\text{scale}}$
Note that the hyperparameter $\lambda_{\text{A}}$ was introduced in our main objective, where it controls the strength of the personalized item alignment regularization. In contrast, $\rho$ and $\lambda_{\text{scale}}$ are two hyperparameters introduced in our algorithm to dynamically adjust $\lambda_{\text{A}}$ during training. Specifically, we use $\rho$ and $\lambda_{\text{scale}}$ to gradually increase the value of $\lambda_{\text{A}}$. We monitor the performance on the validation set to determine whether the latent space becomes trapped in a local optimum i.e., when the performance does not improve after $\rho$ epochs. If such a situation is detected, we increase $\lambda_{\text{A}}$ by multiplying it with the scaling factor $\lambda_{\text{scale}}$.
In this section, we analyze the sensitivity of $\lambda_{\text{A}}$, $\rho$, and $\lambda_{\text{scale}}$ with respect to model performance. We fix the value of $\lambda_{\text{A}}=\{2,4,8\}$ and vary the values of $\rho$ and $\lambda_{\text{scale}}$. Note that $\rho$ is used to detect local optima by checking whether the validation performance fails to improve for $\rho$ consecutive epochs. We evaluate $\rho$ in the range from $3$ to $15$. The hyperparameter $\lambda_{\text{scale}}$ controls the rate at which $\lambda_{\text{A}}$ increases. Since a large scaling factor may destabilize training, we test $\lambda_{\text{scale}}$ values in the range from $1.0$ to $3.0$. $1.0$ mean there will no scaling.
Figure 3: Parameter sensitivity analysis on $\rho$ and $\lambda_{\text{scale}}$ of PIA-VAE on MovieLens-20M. Fix $\lambda_{A}=2.0$ different values of $\rho\in\{3,5,7,10,15\}$ and $\lambda_{\text{scale}}\in\{1.0,1.5,2,0,2.5,3.0\}$
Figure 4: Parameter sensitivity analysis on $\rho$ and $\lambda_{\text{scale}}$ of PIA-VAE on MovieLens-20M. Fix $\lambda_{A}=4.0$ different values of $\rho\in\{3,5,7,10,15\}$ and $\lambda_{\text{scale}}\in\{1.0,1.5,2,0,2.5,3.0\}$
Figure 5: Parameter sensitivity analysis on $\rho$ and $\lambda_{\text{scale}}$ of PIA-VAE on MovieLens-20M. Fix $\lambda_{A}=8.0$ different values of $\rho\in\{3,5,7,10,15\}$ and $\lambda_{\text{scale}}\in\{1.0,1.5,2,0,2.5,3.0\}$
Figures 3, 4, and 5 show the performance measured by nDCG@100 on the MovieLens-20M dataset. It can be observed that the validation nDCG@100 remains relatively stable across $\lambda_{A}$ values, for $\lambda_{\text{scale}}$ values in the range $1.5$ to $2.3$ and for $\rho$ values in the range $5$ to $10$. Based on this analysis, we set $\rho=5$ and $\lambda_{\text{scale}}=2$ for all experimental settings.
Figure 6: Performances on MovieLens-20M dataset with different $\lambda_{\text{A}}$
Analysis on $\lambda_{\text{A}}$
To futher analyze the sensitivity of $\lambda_{\text{A}}$, we fix the values of $\rho=5$ and $\lambda_{\text{scale}}=2$ (based on previous analyses) and vary the value of $\lambda_{\text{A}}$.
As described in the experimental setup, the hyperparameter $\lambda_{\text{A}}$ is selected using nDCG@100 on the validation set as the evaluation metric across all experiments. Figure 6 presents the validation nDCG@100 for different values of $\lambda_{\text{A}}$, along with the corresponding test nDCG@100, Recall@20 and Recall@50 on MovieLens-20M. It can be seen that the validation and test performances are aligned, and the results also indicate that PIA-VAE is generally robust to variations when $\lambda_{\text{A}}>0$.
D.2 Computational Infrastructure and Running Time
Infrastructure
All experiments, including our approach and the baseline models, were conducted on an ml.g5.2xlarge AWS EC2 instance equipped with an NVIDIA A10G GPU.
Table 6: Average training time per epoch (in seconds).
VAE VAE + Alignment
MovieLens-20M 8.207 9.516 (15.95%)
Netflix 22.193 25.722 (15.90%)
Running Time
Table 6 reports the average training time per epoch (in seconds) for the base VAE model and our proposed PIA-VAE. The addition of the alignment component in PIA-VAE results in a training overhead of approximately 15.9% compared to the base VAE.